answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of drawing a queen or a club from a deck of cards?

Probability (P) of A or B is: P(A) + P(B) - P(A and B). Apply to the question is: P(Q) + P(Club) - P (Q&Club) = 4/52 + 13/52 - 1/52 = 16/52 = 4/13 or .3077 or 30.77%.

A vertical line is drawn through a normal distribution at z equals plus 0 25 The line separates the distribution into two sections and the smaller section corresponds to of the whole distribution?

At Z = 0.25, the area (larger) to the left of the line is 0.5987. The smaller area to the right of the line is 1 - 0.5987 = 0.4013. Therefore, the smaller section corresponds to 40.13% of the area.

Given a standard normal distriution find the area under the curve which lies to the left of z equals 1.43?

From the table in the related link, Z = 1.43, area to left is 0.9236. This is read directly from the table.

A pair of dice are rolled find the probability at least one dice is a 4?

When rolling one die, the probability of getting a 4 is 1 in 6, or 0.1667. If two dice are rolled, you get two unrelated chances of rolling at least one 4, so the probability is 2 in 6, or 0.3333.

The percentage that is one standard deviation away from mean?

For normally distributed data. One standard deviation (1σ)

Percentage within this confidence interval
68.2689492% (68.3% )


Percentage outside this confidence interval
31.7310508% (31.7% )

Ratio outside this confidence interval
1 / 3.1514871 (1 / 3.15)

What is probability of drawing a seven or a nine out of cards?

Probability of drawing a 7 or 9 is the probability of draw a 7 plus the probability of draw a 9 which is 4/52 + 4/52 = 8/52 = 2/13 = 0.1538.

The fact that probabilities are proportions means that they?

The fact that probabilities are proportions means that they are less than or equal to 1.

What is the probability of a coin landing on its side?

Prob. of Landing on one side = 0.5
Prob. of Landing on the other = 0.5

Therefore, Prob. of Landing on one side or the other = 0.5 +0.5 = 1

:)





I think you mean 'landing on its edge'. In that case it depends on the coin (some even have rounded edges).


Interestingly what face a coin lands on is not exactly 50/50. Each person has an individual flipping-style which leads to a different side being preferred (not explicitly heads vs. tails but in 'landing with the same face upwards as was upwards at the beginning of the flip' vs. 'landing with the other face upwards to the way it started'). On average (with most people) the coin lands the same way up as it started more often than not (you'll have to do hundreds and hundreds of flips before you notice this).




If (like me) you want to be really pedantic you would also take into account the very slight possibility that the coin won'tland at all.

What is the equation of probability?

The probability of an event occurring can be found by dividing the number of favorable outcomes (what you want to happen) by the number of possible outcomes

number of favorable outcomes

probability = _________________________

number of possible outcomes

How many 5 digit numbers can be formed out of 56789 if they can be repeated?

Five numbers (5, 6, 7, 8, 9) can be selected for each of the 5 digits. So, there are 5*5*5*5*5 or 55 or 3125 five digit numbers.

What is the probability of drawing 3 cards that are all queens from a deck of cards?

Aprox. 0.018%

There are 4 queens in a regular deck of 52 cards.

The probability of drawing a queen on the first draw is: P(Q1) = 4/52.

The probability of drawing a queen on the second draw given that the first card was

a queen is: P(Q2│Q1) = 3/51.

The probability of drawing a queen on the third draw given that the first two cards

were queens is: P(Q3│(Q2UQ1)) = 2/50.

The probability of drawing 3 queens on the first 3 cards drawn from a deck of cards

is: P(Q1UQ2UQ3) = (4/52)∙(3/51)∙(2/50) = 1.80995... x 10-4 ≈ 0.00018 ≈ 0.018%

There are 40 yellow marbles and blue marbles in a bag there are four times as many blue marbles as yellow marbles how many yellow and blue marbles are there?

Set up the problem as: y=number of yellow marbles & 4y=number of blue marbles. 40=y+4y or 40=5y or y=8. Therefore number of blue marbles is 4*8 = 32.

What does one standard deviation mean?

Standard deviation is a measure of variation from the mean of a data set. 1 standard deviation from the mean (which is usually + and - from mean) contains 68% of the data.

What is the probability of obtaining 4 ones in a row when rolling a fair six sided die?

The following is the probability of obtaining 4 ones IN THE FIRST FOUR rolls of a fair die.

Pr(4 1's) = Pr(1)*Pr(1)*Pr(1)*Pr(1) since the events are independent.

Pr(4 1's) = Pr(1)4 = (1/6)4 = 1/1296 = 0.000772

How do you find out a quiz score?

You find out a quiz score by waiting for the quiz to be graded and returned to you. If it is not returned, you ask the teacher or proctor of the quiz.